OpenStudy (anonymous):

subtract and simplify. 3/4 - 11/20

OpenStudy (anonymous):

3/4 - 11/20 15/20-11/20 4/20 =1/5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you help me with another one 2/3 - 1/2? please

OpenStudy (anonymous):

Can i get a medal? And sure..

OpenStudy (anonymous):

2/3 - 1/2 4/6-3/6 =1/6

OpenStudy (anonymous):

sure and thank you

OpenStudy (anonymous):

Anymore??

OpenStudy (anonymous):

yeah lol 3/10 -1/4

OpenStudy (anonymous):

You know how to do it, or you want me to explain??

OpenStudy (anonymous):

3/10 -1/4 6/20-5/20 =1/20

OpenStudy (anonymous):

if you want you can explain the ones i need help with are the ones i got wrong i have 3 more

OpenStudy (anonymous):

WHich are the ones you got wrong??

OpenStudy (anonymous):

all the ones on here i asked for help on but i understand the way you do it

OpenStudy (anonymous):

SO i dont need need to explain?

OpenStudy (anonymous):

no not really but here is one that i really dont understand....A parent died and left an estate to four children. One inherited 2/9 of the estate, the second inherited 19/81, and the thrid inherited 12/27. How much did the fourth inherit? (type a simplified fraction)